# Correlation studies of the 7He excited states

### v2

@@ IV PWBA model for the 2H(6He,1H)7He reaction
Important qualitative features of the 2H(6He,1H)7He reaction can be understood based on the the plane-wave Born approximation (PWBA) model, which well reflects the “quasifree scattering” aspect of the process. Within the historical terminology for this class of reactions we consider the system of three particles, which, in the anti-lab system, are “target” (6He), and composite “beam” (2H), consisting of “spectator” (proton), “participant” (neutron), see Fig. 5. The experiment is performed in the inverse kinematics and we keep “beam” and “target” notation with quotation marks for this Section to avoid confusion here. Completely analytical treatment of the model is possible under two essential physical assumptions:
@@
The partial components of the elastic -matrix and transfer formfactor for the simple -wave only deuteron WF with radial component are defined as
@@ VI Center-of-mass angular distributions
The disadvantage of the small angular range is that the cross section angular distribution is available in the narrow angular range and do not allow detailed comparison with reaction theory calculations to substantiate specific spin-parity prescription. Nevertheless, we are going to demonstrate that important conclusions are possible on the basis of the c.m.s. angular distributions derived in this experiment.
The experimental c.m.s. angular distributions for different excitation energy ranges of 7He are shown in Fig. 11. The data are corrected for registration efficiency of the setup via the MC procedure. The 7He ground state population cross section is found to be somewhat smaller than in Ref. [4]. The obtained angular distributions can be considered as independent of energy within the uncertainty of data.
@@
It can be seen in Fig. 12 (a) that both the 7He g.s. angular distributions of this work and of Ref. [3] are well reproduced with one parameter set fm and fm. It should be noted that spectroscopic factor for the 7He g.s. was taken the same in both calculations, while in the DWBA studies of [4] somewhat different values and were extracted for the 29 MeV data and for the 11.5 MeV data of Ref. [3].
The stability of the angular distributions, demonstrated in Fig. 11, is quite natural if the obtained data is strongly dominated by the same angular momentum transfer ( g.s., and excited states), in the whole excitation energy range MeV available in the experiment. The observation of the low-lying resonance in 7He was declared in [5], which should also exhibit itself in the angular distributions. Unfortunately, it is known, that specifically for the reactions the cross sections have highly analogous profiles. It can be seen in Fig. 12 (b) that to reliably distinguish the and contributions we should go to and need a data with a high statistical confidence. So, one of the ideas, how to prove or disprove the statement about the resonant state in 7He is to perform a dedicated experiment, in which a broad c.m.s. angular range is accessible.
@@ VII.1.2 Suppression of above MeV
Another possible explanation of the experimental asymmetry function, which naturally stems from Eq. (18) is connected with modification of the high-energy tail of the resonance. To check such possibility we have multiplied the population probabilities by Fermi function with energy parameter 1.3 MeV and width parameter 0.15 MeV. The results of calculations are shown in Fig. 17. There is good qualitative agreement at once and also a quantitative agreement with the data may be achieved by small variations of the phase shifts
@@ VII.1.3 The resonant state in 7He
Thus the interpretation proposed in [5] is “phenomenologically possible” (as off-shell T-matrix parameterization describing the data) but can not be unambiguously related to “real physics” of the 7He continuum (described by the on-shell T-matrix).
@@ IX Conclusions
The population cross section for the 7He g.s. is found to be consistent with the data [3] obtained at 11.5 MeV. The c.m. angular distributions for the reaction are practically independent on the 7He excitation energy, supporting the same reaction mechanism and the same dominant up to MeV.
@@
In this work it was demonstrated that the discussed correlation method has the potential to solve the listed burning questions of the 7He continuum, however it is clear that prospective experiments should have higher statistics, higher energy resolution and unambiguous identification of decay channels.

### v3

@@ IV PWBA model for the 2H(6He,1H)7He reaction
Important qualitative features of the 2H(6He,1H)7He reaction can be understood based on the plane-wave Born approximation (PWBA) model, which well reflects the “quasifree scattering” aspect of the process. Within the historical terminology for this class of reactions we consider the system of three particles, which, in the anti-lab system, are “target” (6He), and composite “beam” (2H), consisting of “spectator” (proton), “participant” (neutron), see Fig. 5. The experiment is performed in the inverse kinematics and we keep “beam” and “target” notation with quotation marks for this Section to avoid confusion here. Completely analytical treatment of the model is possible under two essential physical assumptions:
@@
The partial components of the elastic -matrix and transfer formfactor for the simple -wave deuteron WF with radial component are defined as
@@ VI Center-of-mass angular distributions
The disadvantage of the small angular range is that we are unable to make a detailed comparison of the cross section angular distribution with reaction theory calculations to justify a specific spin-parity prescription. Nevertheless, we are going to demonstrate that important conclusions are possible on the basis of the c.m.s. angular distributions derived in this experiment.
The experimental c.m.s. angular distributions for different excitation energy ranges of 7He are shown in Fig. 11. The data are corrected for registration efficiency of the setup via the MC procedure. The obtained angular distributions can be considered as independent of energy within the uncertainty of data, that may be a signal, that the same set of weights of different values is valid.
@@
It can be seen in Fig. 12 (a) that both the 7He g.s. angular distributions of this work and of Ref. [3] can be reproduced only by different parameter sets. Similary, in the DWBA studies of [4] somewhat different spectroscopic factor values and were extracted for the 29 MeV data and for the 11.5 MeV data of Ref. [3]. Moreover it is demonstrated in Fig. 12 (b), that no monotonic function exists, which can describe both data sets. So, the datasets can not be reconciled within typicaly used direct reaction theoretical frameworks, and this is a good indication, that absolute calibration of one of the data sets is actually wrong. This situation calls for a more precise experiment.
The stability of the angular distributions, demonstrated in Fig. 11, is quite natural if the obtained data is strongly dominated by the same angular momentum transfer ( g.s., and excited states), in the whole excitation energy range MeV available in the experiment. The observation of the low-lying resonance in 7He was declared in [5], which should also exhibit itself in the angular distributions. Unfortunately, it is known, that specifically for the reactions the cross sections have highly analogous profiles. It can be seen in Fig. 12 (c) that to reliably distinguish the and contributions we should go to and need a data with a high statistical confidence. So, one of the ideas, how to prove or disprove the statement about the resonant state in 7He is to perform a dedicated experiment, in which a broad c.m.s. angular range is accessible.
@@ VII.1.2 Suppression of above MeV
Another possible explanation of the experimental asymmetry function, which naturally stems from Eq. (18) is connected with modification of the high-energy tail of the resonance. To check such possibility we cut off the high energy tail via multiplication the population probabilities by Fermi function with energy parameter 1.3 MeV and width parameter 0.15 MeV. The results of calculations are shown in Fig. 17. There is good qualitative agreement at once and also a quantitative agreement with the data may be achieved by small variations of the phase shifts
@@ VII.1.3 The resonant state in 7He
Thus, the interpretation proposed in [5] is “phenomenologically possible” as off-shell T-matrix parameterization describing the data, but this interpretation can not be unambiguously related to “real physics” of the 7He continuum described by the on-shell T-matrix.
@@ IX Conclusions
The population cross section for the 7He g.s. is found to be inconsistent with the data [3] obtained at 11.5 MeV. The c.m. angular distributions for the reaction are practically independent on the 7He excitation energy, supporting the same reaction mechanism and the same dominant up to MeV.
@@
In this work it was demonstrated that the discussed correlation method has the potential to solve the listed burning questions of the 7He continuum. However it is clear, that prospective experiments should have higher statistics, higher energy resolution and unambiguous identification of decay channels.